| ID | 
Date | 
Author | 
Group | 
Subject | 
Text | 
  |  
| 
  289  
 | 
05 Aug 2023, 15:33 | 
Ben, Justine | SCS | Shift summary | Achievements
 
 
 
	Measured XAS and RIXS from 
NdNiO3 test sample at 230 K
 
 
 
Plan | 
   |  | 
  288  
 | 
05 Aug 2023, 08:20 | 
Romain Letrun | SPB | Shift summary | X-ray delivery
 
 
 
	9.3 keV, 1-202 pulses at 
1.13 MHz, 1.9-2.1 mJ.
 
	Overall | 
   |  | 
  287  
 | 
04 Aug 2023, 23:20 | 
Andreas, Sergii | SCS | late shift summary | achievements:
 
 
 
	hooked up sample cooling 
and reached 20K at sample at typical low | 
   |  | 
  286  
 | 
04 Aug 2023, 07:10 | 
Giuseppe Mercurio, Teguh Asmara | SCS | night shift summary | achievements:
 
 
 
	changed the LEPM to optimize 
resolution
 
	found 140 meV
 
 
 
problems: | 
   |  | 
  285  
 | 
02 Aug 2023, 06:44 | 
Giuseppe Mercurio, Teguh Asmara | SCS | shift summary | achievements:
 
 
 
	beam transport with the 
second half of undulators closed
 
	measured | 
   |  | 
  284  
 | 
31 Jul 2023, 01:42 | 
Sergii, Martin | SCS | Night shift summary 30.07.2023 | Achivements:
 
 
-We measured vertical spot size 
for several VFM settings | 
   |  | 
  283  
 | 
30 Jul 2023, 23:06 | 
Ulrike Boesenberg | MID | Shift summary | 9 and 10keV beam, up to 30 bunches, significant 
drop in intensity in the morning and in the 
evening. | 
   |  | 
  282  
 | 
30 Jul 2023, 18:35 | 
Giuseppe, Andreas | SCS | Morning Shift summary | achievements:
 
 
 
	calibration of exit slits
 
	measurement 
of Cu L3 edge
 
	moved hRIXS to nominal | 
   |  | 
  281  
 | 
30 Jul 2023, 02:02 | 
Ulrike Boesenberg | MID | Shift summary | 9.3 keV, 1-35pulses, stable delivery ~600uJ
 
*General alignment of MID branch to 
the DES | 
   |  | 
  280  
 | 
29 Jul 2023, 23:46 | 
Adam Round | SPB | Shift end | X-ray delivery
 
 
 
	12.4 keV, up tp 352 pulses 
but the intensity drop across the train was | 
   |  | 
  279  
 | 
29 Jul 2023, 17:57 | 
Harald Sinn | PRC | Issue | The hutch search of SASE3, SXP became defective 
around 5am. 
 
 
DESY MPS group replaced a defective | 
   |  | 
  278  
 | 
28 Jul 2023, 15:37 | 
Harald Sinn | PRC | Status | Good news: The 'jitter'-tests was 
sucessful and BKR could identify the jitter 
source related to a wrong software setting | 
   |  | 
  277  
 | 
28 Jul 2023, 14:56 | 
Giuseppe Mercurio, Laurent Mercadier | SCS | end of shift summary | achievements:
 
 
 
	found new M2 bender position 
for focusing using blank
 
	tried | 
   |  | 
  276  
 | 
28 Jul 2023, 14:46 | 
Harald Sinn | PRC | Status | As discussed, there will be a beam interuption 
form 14:45 until probably 15:45 related to 
jitter studies at A1 that the BKR would like | 
   |  | 
  275  
 | 
28 Jul 2023, 06:38 | 
Sergii, Manuel | SCS | Shift Summary 27_07_23 night | Achievements:
 
 
1. Beam in sample chamber, almost 
overlapping with ALAS | 
   |  | 
  274  
 | 
25 Jul 2023, 19:18 | 
Harald Sinn | PRC | Status | After the 12 hour cryo failure last night, 
it was decided to exchange a faulty window 
with some reduced transmission at the | 
   |  | 
  273  
 | 
23 Jul 2023, 17:28 | 
Frederik Wolff-Fabris | PRC | Status | Photon Activities by XPD colleagues 
in SA3 are concluded on Sunday 23.07 at 17:00. 
The photon commissioning activities are concluded | 
   |  | 
  272  
 | 
21 Jul 2023, 20:31 | 
Silja Schmidtchen | XRO | Shift Summary SA2 | Alignment to HED: https://in.xfel.eu/elog/XRO+Beam+Commissioning/205
 
 
SA2 - divergence with various M2 
bender positions  HED - https://in.xfel.eu/elog/XRO+Beam+Commissioning/200 | 
   |  | 
  271  
 | 
21 Jul 2023, 17:22 | 
Kelin Tasca | XRO | Summary of SA1 shift for XO and instruments | PBLM checked:  https://in.xfel.eu/elog/XRO+Beam+Commissioning/203
 
 
Solid attenuators checked: https://in.xfel.eu/elog/XRO+Beam+Commissioning/194 | 
   |  | 
  270  
 | 
21 Jul 2023, 16:54 | 
Antje Trapp | XRO | Beam aligned to HED | Position stored in XRO setpoints - F and 
HED setpoints - A | 
    |   
 |